OpenStudy (anonymous):

Trick question...will give medal...2+2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

||||=4

OpenStudy (anonymous):

3

OpenStudy (anonymous):

5

sam (.sam.):

100

OpenStudy (anonymous):

I said trick question claw...

OpenStudy (anonymous):

George Orwell's Nineteen Eighty-Fou

OpenStudy (anonymous):

i say 3

OpenStudy (anonymous):

Lmao what?

OpenStudy (anonymous):

Keep guessing!

sam (.sam.):

0=0 1=1 2=10 3=11 4=100

OpenStudy (anonymous):

8

OpenStudy (anonymous):

16

OpenStudy (anonymous):

22

OpenStudy (anonymous):

IT is 5

OpenStudy (anonymous):

*yawn*

OpenStudy (anonymous):

Take any two ordinary constants. (e.g. x and y) Make them alike, as in x = y = any non-zero number (in this case, 2) Multiply each by y, but x works equally well (y²=xy). Subtract x² from each equation (y²-x² = xy-x²). Factor y²-x² into (y+x)(y-x) to make ((y-x)(y+x) = xy-x²). Factor xy-x² into x(y-x) to make ((y-x)(y+x) = x(y-x)). Divide by (y-x) to get (y+x=x). Subtract x from both sides of this equation to get y=0. This means that y=0. but since, y equaled 2 in the first case, that means 2 = 0. If you multiply this equation by 2.5, you get 5 = 0. Since 0 + 0 = 0, you can plug in 2 for the first two zeroes, and five for the last one, and then, TA DA! 2 + 2 = 5.

OpenStudy (anonymous):

In boolean logic, 2+2 is true.

OpenStudy (anonymous):

22

sam (.sam.):

Fish?

OpenStudy (anonymous):

Also in boolean logic, 2+2 is 2

OpenStudy (anonymous):

SAM! YOU WIN!

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

2+2 = 2(1+1) = 2((1+1)^2 - 2*1*1)) = after that... try it your self :P :P

OpenStudy (anonymous):

I am literally laughing my head off!!!!!!

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

people watching to many cartoons

OpenStudy (anonymous):

2+2..reverse one of the two's join it with the other u get a fish shape.

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

Exactly lucky

OpenStudy (anonymous):

Not on my system. I end up with a heart sitting on a line.

OpenStudy (anonymous):

I ain't got no steenking feeshes!

OpenStudy (anonymous):